OpenStudy (anonymous):

Solve by simplifying the problem. There are 18 students in Mr. Alvarez’s art class. Throughout the year, every student must pair up with every other student to complete a project. How many projects will be finished? A. 149 projects B. 153 projects C. 171 projects D. 135 projects

OpenStudy (anonymous):

answer will be 153

OpenStudy (mathstudent55):

17 + 16 + ... + 2 + 1 = ?

OpenStudy (anonymous):

18 student and they each will do project with each other ie in combination of 2 people so We will go for combination 18 c 2 \[18! / (18-2)! \times 2!\]

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

17 + 16 + ... + 2 + 1 = \(\dfrac{17 \cdot 18}{2} = 153\)